A projectile is fired at the top of a building at an angle of 20° below the horizontal. If the muzzle velocity of the projectile is 8 m/s, how high is the building if the projectile reaches the ground after 3sec. 


1
Expert's answer
2021-06-15T10:38:01-0400

The equation to solve for time of the freight is a quadratic equation and is given as follows for object thrown at an angle below the horizontal.

"0.5gT^2+(usin\theta) T-h=0"


u, T and angle are given in the question. The same are substituted to obtain height of the building.


"h= 0.5gT^2+(usin\theta) T\newline
h=0.5\times9.8\times3^2+8\times sin20 \times3\newline

h= 52.31m"